A pharmacist wishes to mix a solution that is 2% minoxidil. She has on hand 50 mL of a 1% solution, and she wishes to add some 4% solution to it to obtain the desired 2% solution. How much of the 4% solution should she add?
The equation that I got .5+.04(50+x)=.2(100+x)
I keep on getting a negative number.

Answer by Paul(988)

1(50)+4x=2(x+50)
50+4x=2x+100
2x=50
x=25

Hence, she should add about 25 mL to the solution of 4%.
Paul.